OpenStudy (anonymous):

4t/t^2-81 + t/t-9 add and simplify if possible

OpenStudy (anonymous):

you have to be more clear with asking these questions. use brackets or parentheses where necessary

OpenStudy (anonymous):

i typed it as written

OpenStudy (anonymous):

(t(t+13))/(t^2-81)

jhonyy9 (jhonyy9):

4t/(t*2 - 81 +t/(t-9) =4t/(t+9)(t-9) + t/(t-9) =(4t+t(t+9))/(t*2 -81) = =(4t+t*2 +9t)/(t*2 -9*2) =(t*2 +13t)/(t*2 -81) =t(t+13)/(t*2 -81)